# Double Configuration Interaction Singles: Scalable and size-intensive approach for orbital relaxation in excited states and bond-dissociation

### v1

@@ Abstract
We present a novel theoretical scheme for orbital optimization in configuration interaction singles (CIS) based on a perturbative treatment of its electronic Hessian, whose analytical derivation is also established in this work. The proposed method, which can be interpreted as a “CIS-then-CIS” scheme, variationally accounts for orbital relaxation in excited states, thus significantly reducing the overestimation of charge-transfer excitation energies commonly associated with standard CIS. Additionally, by incorporating de-excitation effects from CIS, we demonstrate that our approach effectively describes single bond dissociation. Notably, all these improvements are achieved at a mean-field cost, with the pre-factor further reduced with the efficient algorithm introduced here, while preserving the size-intensive property of CIS.
@@ I Introduction
Configuration Interaction Singles (CIS) is the most basic approach to studying excited states, and has been extensively investigated because of its simplicity. It is now well known that CIS tends to overestimate charge transfer excitations, for the reasons mentioned above.Subotnik (2011); Liu et al. (2012) To address this issue, several studies have focused on both approximate and rigorous orbital optimizations for CIS.Liu et al. (2012); Shea and Neuscamman (2018); Shea et al. (2020); Hardikar and Neuscamman (2020) Liu and co-workers perturbatively estimated the first-order orbital rotation effect using the HF Hessian as an approximation to the CIS Hessian.Liu et al. (2012) Shea and Neuscamman applied the generalized variational condition, approximately minimizing the expectation value of the square of the energy-shifted Hamiltonian within the CIS framework.Shea and Neuscamman (2018); Shea et al. (2020) They later performed full orbital optimization by requiring the CIS orbital gradient to vanish.Hardikar and Neuscamman (2020)
@@ II.1 Orbital optimization in CIS
and minimize its energy $E[{\bm{\kappa}},{\bf c}]$. The first derivatives of the energy with respect to $\kappa_{ai}$ and $c_{ai}$ around ${\bm{\kappa}}={\bf 0}$ are
${{}^{\rm o}}g_{ai}$ $=\frac{\partial E[{\bm{\kappa}},{\bf c}]}{\partial\kappa_{ai}}\Big{|}_{{\bm{\kappa}}={\bf 0}}=\langle 0|\left[\hat{E}^{-}_{ai},(\hat{H}-E_{0})\right]|0\rangle$ (3)
@@
The explicit working equations for evaluating these terms are provided in the Supporting Information.si
@@ II.2 Double CIS
Since CIS is invariant with respect to rotations among the occupied and virtual spaces, the parameterization for $\hat{\kappa}$ is sufficient. Hence, an optimized state may be given by
@@
This form is indeed identical to that explored previously by Liu and SubotnikLiu et al. (2012), where $\kappa_{ai}$ were approximated perturbatively with the Hartree-Fock Hessian. They further extended their scheme (OOCIS) to generate several perturbed CIS states and constructed an effective Hamiltonian matrix in this subspace, which was then diagonalized; thus falling into the category of the “perturb-then-diagonalize” approach.Liu et al. (2013); Liu and Subotnik (2014)
@@
Here, $|\bar{\bf c}\rangle=\bar{c}_{0}|\Phi_{0}\rangle+\sum_{ai}\bar{c}_{ai}|\Phi_{i}^{a}\rangle$ represents a generalized CIS ansatz.Shea and Neuscamman (2018) We intentionally introduced the HF determinant $|\Phi_{0}\rangle$, to account for the response in the ground state and its orthogonality. Eq.(14) encompasses the first order corrections to both orbital rotation and CI coefficients in terms of Eq.(1). Nevertheless, it is more convenient to equivalently rearrange the variational parameters, without loss of generality, so that the DCIS state is expressed as
@@
which is regarded as CIS on top of CIS. In fact, it can be easily shownsi that its variational space is identical to that of the “internally-contracted” multi-reference CIS with the CIS reference, i.e., $\sum_{pq}t_{pq}\hat{E}_{pq}|0\rangle$. While the internally-contracted form may be more appealing because of its clear physical interpretation, the form presented in Eq.(15) is preferred in practice for two reasons: (1) the internally-contracted CIS can introduce a significant number of redundancies, posing considerable challenges in optimization whereas Eq.(15) is free of such redundancies, and (2) it allows for a direct translation of the results from the electronic Hessian into the DCIS working equations, requiring minimal modification to the Hessian program, as shown below.
@@
$\langle 0|\hat{E}_{kc}(\hat{H}-E_{\rm DCIS})|\Psi_{\rm DCIS}\rangle$ $=0$ (16)
$\langle\Phi_{\mu}|(\hat{H}-E_{\rm DCIS})|\Psi_{\rm DCIS}\rangle$ $=0$ (17)
Interestingly, the DCIS Hamiltonian matrix, the second derivative of the energy expectation value with respect to $\{d_{ai},\bar{d}_{0},\bar{d}_{ai}\}$, parallels the electronic Hessian matrix H. Namely, the DCIS generalized eigenvalue problem is
@@ II.2.1 Properties of DCIS
Size-intensivity. The excitation energy of DCIS is essentially size-intensive. In other words, for a non-interacting system AB where A is sufficiently distant from B, then the relaxation energy of AB is the same as that computed solely for the system A (or B). However, a caveat arises when the target CIS state is exactly degenerate, like all other state-specific methods. Namely, if A and B are the same molecule having the identical geometry, any linear combination between two “local” CIS states of A and B can form a whole CIS state of AB, which is yet not a product state, i.e., $|\Psi_{\rm CIS}^{\rm AB}\rangle=C_{A}|\Psi_{\rm CIS}^{\rm A}\rangle+C_{B}|\Psi_{\rm CIS}^{\rm B}\rangle\neq|\Psi_{\rm CIS}^{\rm A}\rangle|\Psi_{\rm CIS}^{\rm B}\rangle$. This causes an inconsistency where the DCIS energy is not invariant with respect to rotation between degenerate CIS states. This issue occurs because DCIS is a state-specific, linear method. Fortunately, this is a rare event and we do not expect this to happen in practice. Furthermore, one can easily circumvent this issue by extending DCIS to a multi-state scheme similar to Ref.[Liu et al., 2013], although such an extension is beyond the scope of this work.
Bond dissociation. DCIS can, in principle, describe single bond-breaking as it includes doubly excited configurations $|\Phi_{ij}^{ab}\rangle$ as well as the HF determinant $|\Phi_{0}\rangle$. By diagonalizing the DCIS Hamiltonian, one not only obtains the energy relaxation contribution to the target CIS state, but also other accesses variational states. Of particular interest is the lowest energy solution, which is guaranteed to have a negative excitation energy representing “de-excitation” and a total energy lower than that of the HF state, thus representing the ground state with a large weight $\bar{d}_{0}$. The ground state also benefits from approximate orbital rotation effects, thanks to the presence of singles. It is noteworthy that the coefficients for doubles are constrained since $c_{ai}$ are fixed, and hence its performance must be tested.
@@
Computational scaling. The formal scaling of DCIS remains at $O(N^{4})$ where $I$ is the system size, identical to that of CIS. However, computing the gradients ${{}^{\rm o}}{\bf g}$ requires the formation of two Fock-like matrices. Although this is only necessary once, calculating each of the sigma-vectors with the trial vectors ${}^{\rm o}{\bf x}$ and ${}^{\rm c}{\bf x}$ involves forming two Fock-like matrices for ${{}^{\rm oo}}{\bf A}{{}^{\rm o}}{\bf x}$, and one for each of ${{}^{\rm oc}}{\bf A}{{}^{\rm c}}{\bf x}$, ${{}^{\rm co}}{\bf A}{{}^{\rm o}}{\bf x}$, and ${{}^{\rm cc}}{\bf A}{{}^{\rm c}}{\bf x}$. As a result, DCIS is approximately five times more expensive than CIS per cycle. However, this disadvantage is largely offset by the fact that CIS requires determining basically all the low-lying $(I-1)$ states, whereas DCIS can bypass this computational burden, as discussed below.
@@ II.2.2 Efficient computational scheme
However, given that the target DCIS state $|\Psi_{\rm DCIS}\rangle$ is expected to represent a small (perturbative) correction to the CIS state $|0_{I}\rangle$, we can “follow” the target state during iterative optimization with the Davidson algorithm, bypassing the other low-lying eigenstates. Specifically, we initialize with $|0_{I}\rangle$ (i.e., $\bar{d}_{ai}=c_{ai}^{I}$ and $\bar{d}_{0}=d_{ai}=0$) and compute the overlap between the CIS and trial DCIS states, $|\langle 0_{I}|\Psi^{(k)}_{K,\rm DCIS}\rangle|$, at each $k$th cycle, to identify the corresponding trial vector $|\Psi^{(k)}_{I,\rm DCIS}\rangle$ with the largest overlap. We then form the perturbative update only for $|\Psi^{(k)}_{I,\rm DCIS}\rangle$ by using diagonal preconditioning, as opposed to the block Davidson algorithm,Sadkane and Sidje (1999) which updates multiple states simultaneously and therefore requires to calculating Fock-like matrices many times. While alternative methods like DIISPulay (1980, 1982); Hardikar and Neuscamman (2020) could be employed, the proposed "overlap-following" algorithm guarantees orthogonality from other low-lying states and is stable, with a simple modification to the Davidson algorithm.
@@ III Results
We have implemented both the CIS Hessian and DCIS in the Gellan suite of programs. However, it is not our intention to investigate the performance of Hessian-based orbital optimization of CIS; its detailed investigation will be conducted in future studies. Instead, we focus here on DCIS and demonstrate its capabilities. To do so, we first examine its size-intensive properties, and present results for charge transfer excitations and bond dissociation.
@@ III.1 Size-intensivity test
As direct numerical evidence of the size-intensivity of DCIS, we calculated the DCIS relaxation energy of a CIS excitation using the HF molecule with $n$ Be atoms separated by 10 Å. Table 1 shows the CIS excitation energy of $\pi-\pi^{*}$ transition computed at a bond length of $1.0$ using a 6-31G basis set for $n=0,1,2$, along with the DCIS relaxation effect on this CIS state. Clearly, the results confirm that DCIS is size-intensive, as the relaxation remains unchanged with system size. Additionally, the DCIS ground state, as the lowest root, exhibits the same invariance character, demonstrating that the correlation energy is size-intensive, though not size-consistent, similar to spin-flip (SF) CIS.Krylov (2001)
@@ III.2 Charge transfer excitations
For charge transfer excitations, the orbital derivative ${}^{\rm o}{\bf g}$ is large, indicating that a significant orbital relaxation effect is needed to lower the energies and mitigate the HF orbital bias. Since DCIS incorporates this relaxation in CIS, it is expected to improve excitation energies, particularly compared to OOCIS, because the latter approximates the CIS Hessian by that of HF.Liu et al. (2012)
@@
To demonstrate this, we computed the CT excitation energies for 30 states across 17 organic compounds (19 structures), using the data set from Loos et al..Loos et al. (2021) These results were compared with the theoretical best estimates based on coupled cluster methods with cc-pVTZ.Loos et al. (2021) Fig.1 summarizes the errors in excitation energies for these CT states, ordered in ascending order of the CIS gradient $|^{\rm o}{\bf g}|$ (red line). As expected, the CIS error generally increases with the magnitude of $|^{\rm o}{\bf g}|$, while DCIS significantly improves the excitation energies. Note that there are two reference ground state options for the DCIS results. If the HF state is used as the ground state, as in OOCIS, then the results directly reflect the relaxation of the excited state. Alternatively, one can use the DCIS lowest eigenstate, which preserves orthogonality. While the former choice (DCIS/HF) usually outperforms the latter (DCIS/DCIS), the states inevitably become nonorthogonal.
@@
Fig. 2 presents the relaxation energy provided by OOCIS and DCIS for each excited state as a function of the CIS gradient. As expected, the OOCIS relaxation effect is consistently weaker than that of DCIS, which rigorously employs part of the CIS Hessian information (A). Both methods yield similar results for states with $|^{\rm o}{\bf g}|<0.15$ but the difference becomes more pronounced with increasing $|^{\rm o}{\bf g}|$. Overall, we find that DCIS is a better choice compared to OOCIS, particularly for states with significant CT character, as indicated by their mean-errors (MEs) and mean-absolute-errors (MAEs) in Table 2. The OOCIS results for systems with $|^{\rm o}{\bf g}|>0.15$ deteriorate by approximately 0.2 eV on average, whereas DCIS maintains more or less similar accuracy, independent of $|^{\rm o}{\bf g}|$. In passing, the DCIS ground state correlation energy is sufficiently small across all systems as shown in Fig. 2, further supporting its size-intensivity. However, the magnitude of the ground state correlation energy is typically falls within the range of 0.1-0.2 eV, and therefore the DCIS/DCIS results are slightly overestimated compared to those of DCIS/HF.
Table 2: ME and MAE of each method.
@@
    Systems with $|^{\rm o}{\bf g}|>0.15$.
Figure 3: Potential energy curves for (a) HF and (b) F2. Errors from FCI are depicted in (c) and (d). NPEs of RHF, DCIS, and SFCIS are 247.2, 14.7, and 81.7 mH for the HF molecule, and 342.8, 15.9, and 38.4 mH for the F2 molecule.

### v3

@@ Abstract
We present a novel theoretical scheme for orbital relaxation in configuration interaction singles (CIS) based on a perturbative treatment of its electronic Hessian, whose analytical derivation is also established in this work. The proposed method, which can be interpreted as a “CIS-then-CIS” scheme, variationally accounts for orbital relaxation in excited states, thus significantly reducing the overestimation of charge-transfer excitation energies commonly associated with standard CIS. Additionally, by incorporating de-excitation effects from CIS, we demonstrate that our approach effectively describes single bond dissociation. Notably, all these improvements are achieved at a mean-field cost, with the pre-factor further reduced with the efficient algorithm introduced here, while preserving the size-intensive property of CIS.
@@ I Introduction
Configuration Interaction Singles (CIS) is the most basic approach to studying excited states, and has been extensively investigated because of its simplicity. It is now well known that CIS tends to overestimate charge transfer excitations, for the reasons mentioned above.Subotnik (2011); Liu et al. (2012) To address this issue, several studies have focused on both approximate and rigorous orbital optimizations for CIS.Liu et al. (2012); Shea and Neuscamman (2018); Shea et al. (2020); Hardikar and Neuscamman (2020) Liu and co-workers perturbatively estimated the orbital rotation effect by truncating the CIS energy as a quadratic function of orbital rotation parameters, which is then minimized by solving a linear equation as a Newton-Raphson step. They further approximated the second derivative of the CIS energy by the HF Hessian.Liu et al. (2012) However, it was found that the correction to charge transfer states by this treatment was generally insufficient.Liu et al. (2013) Shea and Neuscamman applied the generalized variational condition, approximately minimizing the expectation value of the square of the energy-shifted Hamiltonian within the CIS framework.Shea and Neuscamman (2018); Shea et al. (2020) They later performed full orbital optimization by requiring the CIS orbital gradient to vanish.Hardikar and Neuscamman (2020)
@@ II.1 Orbital optimization in CIS
and $\hat{E}_{ai}=a_{a}^{\dagger}a_{i}$, and minimize its energy $E[{\bm{\kappa}},{\bf c}]$. The first derivatives of the energy with respect to $\kappa_{ai}$ and $c_{ai}$ around ${\bm{\kappa}}={\bf 0}$ are
${{}^{\rm o}}g_{ai}$ $=\frac{\partial E[{\bm{\kappa}},{\bf c}]}{\partial\kappa_{ai}}\Big{|}_{{\bm{\kappa}}={\bf 0}}=\langle 0|\left[(\hat{E}_{ai}-\hat{E}_{ia}),(\hat{H}-E_{0})\right]|0\rangle$ (3)
@@
The explicit working equations for evaluating these terms are provided in the Supplementary Material.si
@@ II.2 Double CIS
Since CIS is invariant with respect to rotations among the occupied and virtual spaces, the parameterization for $\hat{\kappa}$ as given in Eq.(2) is sufficient. Hence, an optimized state may be given by
@@
This form is indeed identical to that explored previously by Liu and SubotnikLiu et al. (2012), called orbital-optimized CIS (OOCIS), where $\kappa_{ai}$ were approximated perturbatively with the Hartree-Fock Hessian ${{}^{\rm oo}}{\bf H}^{\rm HF}$. Specifically, they considered the following quadratic energy function,
$E_{\rm OOCIS}[{\bm{\kappa}}]=E_{0}+\sum_{ai}{{}^{\rm o}}g_{ai}\kappa_{ai}+\frac{1}{2}\sum_{abij}{{}^{\rm oo}}H^{\rm HF}_{ai,bj}\kappa_{ai}\kappa_{bj}$ (14)
and minimized it by solving the linear equation
$\sum_{bj}{{}^{\rm oo}}H^{\rm HF}_{ai,bj}\kappa_{bj}+{{}^{\rm o}}g_{ai}=0$ (15)
They further extended their scheme to generate several perturbed CIS states and constructed an effective Hamiltonian matrix in this subspace, which was then diagonalized; thus falling into the category of the “perturb-then-diagonalize” approach.Liu et al. (2013); Liu and Subotnik (2014)
@@
Here, $|\bar{\bf c}\rangle=\bar{c}_{0}|\Phi_{0}\rangle+\sum_{ai}\bar{c}_{ai}|\Phi_{i}^{a}\rangle$ represents a generalized CIS ansatz.Shea and Neuscamman (2018) We intentionally introduced the HF determinant $|\Phi_{0}\rangle$, to account for the response in the ground state and its orthogonality. Eq.(16) clearly encompasses the first order corrections to both orbital rotation and CI coefficients in terms of Eq.(1). It should be emphasized that our method differs from the OOCIS scheme by having a well-defined wave function (16), whose energy is given as the expectation value of the Hamiltonian, whereas there is no associated wave function in OOCIS.
It is more convenient to equivalently rearrange the variational parameters in Eq.(16), without loss of generality, so that the DCIS state is expressed as
@@
which is regarded as CIS on top of CIS. In fact, it can be easily shownsi that its variational space is identical to that of the “internally-contracted” multi-reference CIS with the CIS reference, i.e., $\sum_{pq}t_{pq}\hat{E}_{pq}|0\rangle$. While the internally-contracted form may be more appealing because of its clear physical interpretation, the form presented in Eq.(17) is far preferred in practice for two reasons: (1) the internally-contracted CIS can introduce a significant number of redundancies, posing considerable challenges in optimization whereas Eq.(17) is free of such redundancies, and (2) it allows for a direct translation of the results from the electronic Hessian into the DCIS working equations, requiring minimal modification to the Hessian program, as shown below.
@@
$\langle 0|\hat{E}_{kc}(\hat{H}-E_{\rm DCIS})|\Psi_{\rm DCIS}\rangle$ $=0$ (18a)
$\langle\Phi_{\mu}|(\hat{H}-E_{\rm DCIS})|\Psi_{\rm DCIS}\rangle$ $=0$ (18b)
Interestingly, the DCIS Hamiltonian matrix, the second derivative of the energy expectation value with respect to $\{d_{ai},\bar{d}_{0},\bar{d}_{ai}\}$, parallels the electronic Hessian matrix H. Namely, these equations are recast as the DCIS generalized eigenvalue problem:
@@ II.2.1 Properties of DCIS
Size-intensivity. The excitation energy of DCIS is essentially size-intensive. In other words, for a non-interacting system AB where A is sufficiently distant from B, the relaxation energy of AB is the same as that computed solely for the system A (or B). However, a caveat arises when the target CIS state is exactly degenerate, like all other state-specific methods. Namely, if A and B are the same molecule having the identical geometry, any linear combination between two “local” CIS states of A and B can form a whole CIS state of AB, which is yet not a product state, i.e., $|\Psi_{\rm CIS}^{\rm AB}\rangle=C_{A}|\Psi_{\rm CIS}^{\rm A}\rangle|\Phi_{0}^{\rm B}\rangle+C_{B}|\Phi_{0}^{\rm A}\rangle|\Psi_{\rm CIS}^{\rm B}\rangle\neq|\Psi_{\rm CIS}^{\rm A}\rangle|\Psi_{\rm CIS}^{\rm B}\rangle$. This causes an inconsistency where the DCIS energy is not invariant with respect to rotation between degenerate CIS states. This issue occurs because DCIS is a state-specific, linear method. Fortunately, this is a rare event and we do not expect this to happen in practice. Furthermore, one can easily circumvent this issue by extending DCIS to a multi-state scheme similar to Ref.[Liu et al., 2013], although such an extension is beyond the scope of this work.
Bond dissociation. DCIS can, in principle, describe single bond-breaking as it includes doubly excited configurations $|\Phi_{ij}^{ab}\rangle$ as well as the HF determinant $|\Phi_{0}\rangle$. By diagonalizing the DCIS Hamiltonian, one not only obtains the energy relaxation contribution to the target CIS state, but also accesses other variational states. Of particular interest is the lowest energy solution, which is guaranteed to have a negative excitation energy representing “de-excitation” and a total energy lower than that of the HF state, thus representing the ground state with a large weight $\bar{d}_{0}$. The ground state also benefits from approximate orbital rotation effects, thanks to the presence of singles. It is noteworthy that the coefficients for doubles are constrained since $c_{ai}$ are fixed, and hence its performance must be tested.
@@
Computational scaling. The formal scaling of DCIS remains at $O(N^{4})$ where $N$ is the system size, identical to that of CIS. However, computing the gradients ${{}^{\rm o}}{\bf g}$ requires the formation of two Fock-like matrices.si Although this is only necessary once, calculating each of the sigma-vectors with the trial vectors ${}^{\rm o}{\bf x}$ and ${}^{\rm c}{\bf x}$ involves forming five Fock-like matrices to evaluate ${{}^{\rm oo}}{\bf A}{{}^{\rm o}}{\bf x}$, ${{}^{\rm oc}}{\bf A}{{}^{\rm c}}{\bf x}$, ${{}^{\rm co}}{\bf A}{{}^{\rm o}}{\bf x}$, and ${{}^{\rm cc}}{\bf A}{{}^{\rm c}}{\bf x}$.si As a result, DCIS is approximately five times more expensive than CIS per cycle. However, this disadvantage is largely offset by the fact that CIS requires determining basically all the low-lying $(I-1)$ states, whereas DCIS can bypass this computational burden, as discussed below.
@@ II.2.2 Efficient computational scheme
However, given that the target DCIS state $|\Psi_{\rm DCIS}\rangle$ is expected to represent a small (perturbative) correction to the CIS state $|0_{I}\rangle$, we can “follow” the target state during iterative optimization with the Davidson algorithm, bypassing the other low-lying eigenstates. Specifically, we initialize with $|0_{I}\rangle$ (i.e., $\bar{d}_{ai}=c_{ai}^{I}$ and $\bar{d}_{0}=d_{ai}=0$) and compute the overlap between the CIS and trial DCIS states, $|\langle 0_{I}|\Psi^{(k)}_{K,\rm DCIS}\rangle|$, at each $k$th cycle, to identify the corresponding trial vector $|\Psi^{(k)}_{I,\rm DCIS}\rangle$ with the largest overlap. We then form the perturbative update only for $|\Psi^{(k)}_{I,\rm DCIS}\rangle$ by using diagonal preconditioning, as opposed to the block Davidson algorithm,Sadkane and Sidje (1999) which updates multiple states simultaneously and therefore requires to calculate Fock-like matrices many times. While alternative methods like DIISPulay (1980, 1982); Hardikar and Neuscamman (2020) could be employed, the proposed "maximum-overlap" algorithm guarantees orthogonality from other low-lying states and is stable, with a simple modification to the Davidson algorithm. The pseudocode is presented in the Supplementary Material.si
We note that the maximum-overlap Davidson algorithm does not suffer from the root-flipping issue commonly encountered in its SCF variants.Gilbert et al. (2008); Hait and Head-Gordon (2020b); Carter-Fenk and Herbert (2020) Even if the target trial state $|\Psi^{(k)}_{I,\rm DCIS}\rangle$ exhibits some degeneracy, such that the overlaps $|\langle 0_{I}|\Psi^{(k)}_{I,\rm DCIS}\rangle|$ and $|\langle 0_{I}|\Psi^{(k)}_{J\neq I,\rm DCIS}\rangle|$ accidentally become identical throughout the calculation, convergence is still assured. This is because, at each iteration, a new trial state orthogonal to the current subspace is always added, ensuring that the Krylov subspace continues to expand and eventually spans the entire DCIS Hilbert space. In such cases, the algorithm may converge more slowly, potentially requiring more Fock builds than the block Davidson algorithm. However, we have not encountered this issue in practice. As demonstrated below, the maximum-overlap algorithm achieves significant speed-ups compared to the block Davidson algorithm.
@@ III Results
We have implemented both the CIS Hessian and DCIS in our in-house code. However, it is not our intention to investigate the performance of Hessian-based orbital optimization of CIS; its detailed investigation will be conducted in future studies. Instead, we focus here on DCIS and demonstrate its capabilities. To do so, we first examine its size-intensive property, and present results for charge transfer excitations and bond dissociation. We will then demonstrate that the maximum-overlap algorithm considerably reduces the computational cost compared to the standard Davidson algorithm.
@@ III.1 Size-intensivity test
As direct numerical evidence of the size-intensivity of DCIS, we calculated the DCIS relaxation energy of a CIS excitation using the HF molecule with $n$ Be atoms separated by 10 Å. Table 1 shows the CIS excitation energy of $\pi-\sigma^{*}$ transition computed at a bond length of $1.0$ using a 6-31G basis set for $n=0,1,2$, along with the DCIS relaxation effect on this CIS state. Clearly, the results confirm that DCIS is size-intensive, as the relaxation remains unchanged with system size. Additionally, the DCIS ground state, as the lowest root, exhibits the same invariance character, demonstrating that the correlation energy is size-intensive, though not size-consistent, similar to spin-flip (SF) CIS.Krylov (2001)
@@ III.2 Charge transfer excitations
For charge transfer excitations, the orbital derivative ${}^{\rm o}{\bf g}$ is typically large, indicating that a significant orbital relaxation effect is needed to lower the energies and mitigate the HF orbital bias. Since DCIS incorporates this relaxation in CIS, it is expected to improve excitation energies, particularly compared to OOCIS, because the latter approximates the CIS Hessian by that of HF.Liu et al. (2012)
@@
To demonstrate this, we computed the charge transfer excitation energies for 30 states across 17 organic compounds (19 structures), using the data set from Loos et al..Loos et al. (2021) These results were compared with the theoretical best estimates based on coupled cluster methods with cc-pVTZ.Loos et al. (2021) Fig.1 summarizes the errors in excitation energies for these charge transfer states, ordered in ascending order of the CIS gradient $|^{\rm o}{\bf g}|$ (red line). As expected, the CIS error generally increases with the magnitude of $|^{\rm o}{\bf g}|$, while DCIS significantly improves the excitation energies. Note that there are two reference ground state options for the DCIS results. If the HF state is used as the ground state, as in OOCIS, then the results directly reflect the relaxation of the excited state. Alternatively, one can use the DCIS lowest eigenstate, which preserves orthogonality and facilitates straightforward evaluations of inter-state couplings such as transition dipole moments. While the former choice (DCIS/HF) usually outperforms the latter (DCIS/DCIS), the states inevitably become nonorthogonal.
@@
Fig. 2 presents the relaxation energy provided by OOCIS and DCIS for each excited state as a function of the CIS gradient. As expected, the OOCIS relaxation effect is consistently weaker than that of DCIS, which rigorously employs part of the CIS Hessian information (A). Both methods yield similar results for states with $|^{\rm o}{\bf g}|<0.15$ but the difference becomes more pronounced with increasing $|^{\rm o}{\bf g}|$. In passing, the DCIS ground state correlation energy is sufficiently small across all systems as shown in Fig. 2, further supporting its size-intensivity. However, the magnitude of the ground state correlation energy is typically falls within the range of 0.1-0.2 eV, and therefore the DCIS/DCIS results are slightly overestimated compared to those of DCIS/HF.
While the orbital gradient of CIS energy has a fairly strong correlation with the relaxation energy as shown in Fig. 2 ($R^{2}=0.951$ and $0.955$ for DCIS and OOCIS, respectively), it does not provide direct insight into how much these relaxation energies actually improve the description of excitations involving spatially significant charge transfer. That is to say, the large $|^{\rm o}{\bf g}|$ does not always mean strong charge transfer.
The extent of charge transfer is often characterized by the difference between the dipole moments of the ground and excited states, $|{\bm{\Delta}}{\bm{\mu}}|$. In Fig. 3, we found that $|^{\rm o}{\bf g}|$ is also linearly correlated with $|{\bm{\Delta}}{\bm{\mu}}|$ with $R^{2}=0.745$, though the correlation is weaker than with the relaxation energy. This motivates further investigation into whether DCIS provides a better mean-field description of excited states with significant charge transfer, using $|{\bm{\Delta}}{\bm{\mu}}|$ as a metric. Such a test should also offer more detailed insight into whether DCIS and OOCIS can be useful for charge transfer states.
To that end, in Fig. 4, we plotted the same excitation energy errors shown in Fig. 1 but now as a function of $|{\bm{\Delta}}{\bm{\mu}}|$, as computed by HF and CIS. The results in Fig. 4 clearly demonstrate that DCIS consistently improves the excitation energy for large $|{\bm{\Delta}}{\bm{\mu}}|$ (greater than about 2.5 a.u.) more significantly than OOCIS does. On the other hand, as expected, there is little to no improvement with DCIS or OOCIS over CIS when $|{\bm{\Delta}}{\bm{\mu}}|$ is smaller, with the exception of the dipeptide (8). Although $|{\bm{\Delta}}{\bm{\mu}}|$ is only $0.92$ a.u. for this CIS state, which is arguably too small to be classified as charge transfer, we observed a notable reduction in error with DCIS, by 1.4 eV, due to the large CIS orbital gradient, $|^{\rm o}{\bf g}|\approx 0.2$ a.u.. This large orbital rotation effect and error reduction are thus attributed not to charge transfer, but to the state simply being high-lying, as it corresponds to the 24th CIS state.
We have summarized the mean-errors (MEs) and mean-absolute-errors (MAEs) of each method in Table 2. Our results indicate that DCIS/HF maintains similar accuracy regardless of the excitation character, offering a balanced description. As discussed above, DCIS/DCIS is slightly less accurate than DCIS/HF, mainly due to the small amount of ground state correlation introduced, which leads to an overestimation of valence excitations (systems with $|{\bm{\Delta}}{\bm{\mu}}|<0.25$). Nonetheless, both DCIS/HF and DCIS/DCIS significantly improve the description of charge transfer excitations ($|{\bm{\Delta}}{\bm{\mu}}|>0.25$), achieving MAEs of 0.48 and 0.52 eV, respectively. In stark contrast, OOCIS actually becomes less accurate for charge transfer states, with an MAE of 0.98 eV, which is approximately 0.3 eV worse than for valence states.
Overall, while OOCIS reduces the errors in CIS charge transfer states by half, DCIS achieves an additional reduction of the remaining error by half. However, it is important to note that the orbital rotation effect in DCIS is still an approximation of the full orbital rotation. Further studies are necessary to evaluate the significance of higher-order terms present in the full orbital rotation.
Figure 3: Correlation between CIS orbital gradient ($|^{\rm o}{\bf g}|$) and dipole change ($|{\bm{\Delta}}{\bm{\mu}}|$) using the same test set as Fig. 1.
Figure 4: Error in excitation energy computed with CIS (blue), OOCIS (brown), and DCIS/DCIS (green), as a function of the dipole change. The error area is filled with color for visual guidance.
Table 2: ME and MAE of each method in eV.
@@
    Systems with $|{\bm{\Delta}}{\bm{\mu}}|<2.5$.
  • b
    Systems with $|{\bm{\Delta}}{\bm{\mu}}|>2.5$.
Figure 5: Potential energy curves for (a) HF and (b) F2. Errors from FCI are depicted in (c) and (d).
Table 3: Non-parallelity error (NPE) in mH, defined as the difference between the maximum and minimum errors from the FCI potential cruve.
HF DCIS SFCIS CASSCF(2e,2o)
HF 247.1 14.7 81.7 20.3
F2 342.8 15.9 38.4 24.6
Figure 6: Convergence behavior of block Davidson and maximum-overlap Davidson algorithms. (a) $S_{3}$ of benzonitrile (4). (b) $S_{13}$ of $\beta$-Dipeptide (9). (c) $S_{46}$ of $\beta$-Dipeptide (9).